# Electromagnetic Inverse Scattering from a Single Transmitter

### v5

@@ Abstract
Electromagnetic Inverse Scattering Problems (EISP) seek to reconstruct relative permittivity from scattered fields and are fundamental to applications like medical imaging. This inverse process is inherently ill-posed and highly nonlinear, making it particularly challenging, especially under sparse transmitter setups, e.g., with only one transmitter. While recent machine learning-based approaches have shown promising results, they often rely on time-consuming, case-specific optimization and perform poorly under sparse transmitter setups. To address these limitations, we revisit EISP from a data-driven perspective. The scarcity of transmitters leads to an insufficient amount of measured data, which fails to capture adequate physical information for stable inversion. Accordingly, we propose a fully end-to-end and data-driven framework that predicts the relative permittivity of scatterers from measured fields, leveraging data distribution priors to compensate for the incomplete information from sparse measurements. This design enables data-driven training and feed-forward prediction of relative permittivity while maintaining strong robustness to transmitter sparsity. Extensive experiments show that our method outperforms state-of-the-art approaches in reconstruction accuracy and robustness. Notably, we demonstrate, for the first time, high-quality reconstruction from a single transmitter. This work advances practical electromagnetic imaging by providing a new, cost-effective paradigm to inverse scattering.
Equal contribution, random order.
Correspondence: {ez_cheng, tianchunxun, ou524u}@stu.pku.edu.cn
{wtzhu, maxiaoxuan, yizhou.wang}@pku.edu.cn
Project page: https://gomenei.github.io/SingleTX-EISP/
@@ 1 Introduction
Electromagnetic waves can penetrate object surfaces, making them essential for non-invasive imaging (Geng et al., 2024; O’Loughlin et al., 2018). At the core of electromagnetic imaging lies the Electromagnetic Inverse Scattering Problems (EISP), which seeks to reconstruct an object’s relative permittivity from measured scattered electromagnetic field (Nikolova, 2011). By solving EISP, we can accurately recover internal structures without physical intrusion (Song et al., 2005), enabling a range of scientific and industrial applications, such as safer and more cost-effective alternatives to X-rays and MRI scans (Bevacqua et al., 2021; O’Loughlin et al., 2018; Nikolova, 2011). Typically, EISP necessitate a large number of transmitters and receivers to acquire sufficient measurement data. This requirement, however, leads to increased operational time and higher costs, thereby limiting the practical applicability of electromagnetic imaging techniques (Leigsnering et al., 2011). In contrast, reducing the number of transmitters offers significant advantages, including lower costs and easier deployment in constrained environments.(Baraniuk and Steeghs, 2007; Anitori et al., 2010)
Figure 1: Comparison between our method and the previous state-of-the-art. Left: Img-Interiors (Luo et al., 2024) requires case-specific optimization to reconstruct the permittivity. In contrast, our method is a data-driven framework that operates in an end-to-end, feed-forward manner for solving inverse scattering. Right: Our method yields more accurate reconstructions than Img-Interiors (Luo et al., 2024). It remains robust even with a single transmitter and achieves real-time inference with over $70,000\times$ speed-up.
@@
However, the inherent ill-posed nature of EISP poses significant challenges to accurate reconstruction (Pan et al., 2011; Chen, 2018; Li et al., 2019; Zhong et al., 2016; Luo et al., 2024), particularly when only a limited number of transmitters are available. The scarcity of transmitters leads to an insufficient amount of measured data, which fails to capture adequate physical information for stable inversion. As a result, approaches relying solely on physical mechanisms(Slaney et al., 1984; Belkebir et al., 2005; Chen, 2009; Zhong and Chen, 2011) often fail to achieve accurate reconstruction. Conventional numerical methods such as backpropagation (BP) (Belkebir et al., 2005), generally fail to produce reliable reconstructions under such limited-data conditions. Recent machine learning-based approaches like PGAN (Song et al., 2021) and Physics-Net (Liu et al., 2022) often start with an initial solution derived from numerical methods, i.e., BP, and frame the problem as an image-to-image translation task. With only a limited number of transmitters available, reliance on BP becomes a critical bottleneck. When BP fails, these methods are unable to correct its errors, as they are not fully end-to-end, ultimately leading to inaccurate reconstructions. The most recent method Img-Interiors (Luo et al., 2024) integrates physical mechanisms into neural networks and performs case-by-case optimization. However, in limited-transmitter scenarios, even after optimization has converged, the resulting reconstructions may still diverge substantially from the ground truth (Fig.˜2), underscoring the intrinsic ambiguity of the inverse problem.
@@
To address these limitations, we propose a fully end-to-end and data-driven framework that predicts the relative permittivity of scatterers from measured fields, leveraging data distribution priors to compensate for the scarcity of observational data. Specifically, our model takes the measured fields and the spatial coordinate of a position as input and directly predicts the relative permittivity at that location using Multilayer Perceptron (MLP)s, and is trained in a fully end-to-end manner against the ground-truth data. Our approach bypasses traditional numerical methods like BP, thereby avoiding the inherent constraints associated with conventional inversion techniques in limited-transmitter scenarios and fully exploiting the advantages of data-driven learning. This simple yet effective design enables efficient training across datasets and supports fast, feed-forward inference to achieve accurate and stable reconstruction predictions.
@@ 2 Related Work
2.1 Electromagnetic Inverse Scattering Problems (EISP)
Solving EISP is to determine the relative permittivity of the scatterers based on the scattered field measured by the receivers, thereby obtaining internal imaging of the object. The primary challenges of EISP arise from its nonlinearity, ill-posedness, and errors introduced by the discretization (Pan et al., 2011; Chen, 2018; Li et al., 2019; Zhong et al., 2016; Luo et al., 2024). Traditional methods for solving EISP can be categorized into non-iterative (Slaney et al., 1984; Devaney, 1981; Habashy et al., 1993; Belkebir et al., 2005) and iterative (Chen, 2009; Zhong and Chen, 2011; Xu et al., 2017; Habashy et al., 1994; van den Berg et al., 1999) approaches. Non-iterative methods, such as the Born approximation (Slaney et al., 1984), the Eytov approximation (Devaney, 1981; Habashy et al., 1993), and the BP method (Belkebir et al., 2005), solve nonlinear equations through linear approximations, which inevitably lead to poor quality of the results. For better reconstruction quality, iterative methods (Zhong and Chen, 2009; Chen, 2009; Zhong and Chen, 2011; Xu et al., 2017; Habashy et al., 1994; van den Berg et al., 1999; Gao et al., 2015) such as 2-fold Subspace Optimization Method (SOM) (Zhong and Chen, 2009) and Gs SOM (Chen, 2009) are proposed. To further overcome the ill-posedness of EISP, diverse regularization approaches and prior information have been widely applied (Oliveri et al., 2017; Shen et al., 2014; Liu et al., 2018; Anselmi et al., 2018). However, all of these methods are not generalizable and can be time-consuming because of the iterative schemes (Liu et al., 2022).
@@ 2.2 Machine Learning for EISP
Recent studies shift to leverage neural networks to solve this problem and demonstrate promising results (Geng et al., 2021; Li et al., 2024). Some work (Wei and Chen, 2019; Li et al., 2019; Zhang et al., 2020; Xu et al., 2021; Sanghvi et al., 2019; Song et al., 2021; Liu et al., 2022) adopt a two-stage strategy: they use non-iterative methods such as BP (Belkebir et al., 2005) to generate initial estimates, which are then refined using image-to-image neural networks. While these approaches offer a degree of generalization, they are not end-to-end and remain dependent on BP initialization (Belkebir et al., 2005), which becomes their bottleneck. When physical data are too insufficient to reconstruct the scatterer, especially under single-transmitter settings, these approaches tend to “hallucinate” outputs according to unreliable initialization rather than predict the scatterer based on measured field (see Fig.˜5). A more recent approach, Img-Interiors (Luo et al., 2024), integrates scattering mechanisms into the network architecture and achieves accurate reconstructions. However, it requires case-specific optimization, limiting generalization and making it vulnerable to local minima, often leading to failure in complex settings (see Figs.˜4 and 6). Moreover, it fails under a single transmitter setting even when the optimization may have already converged because of ambiguity. While our method is also learning-based, it is an end-to-end feed-forward framework that simultaneously achieves generalization through data-driven learning. As a result, it consistently outperforms SOTA methods, particularly in the challenging single-transmitter setting where previous approaches fail.
@@ 3 Revisiting EISP
Preliminary. In the forward process, the transmitters produce incident electromagnetic field $\mathbf{E}^{\text{i}}$ to the scatterer, generating scattered electromagnetic field $\mathbf{E}^{\text{s}}$. EISP is the inverse problem of the forward process. That is, for an unknown scatterer, we apply certain incident field $\mathbf{E}^{\text{i}}$ to it, and measure the scattered field $\mathbf{E}^{\text{s}}$ as our input. Our goal is to reconstruct the relative permittivity $\bm{\epsilon}_{r}$ throughout the scatterer. For a detailed background introduction of the physical model of EISP, please refer to our supplementary material (Appx.˜B).
Specifically, the incident field $\mathbf{E}^{\text{i}}$ excites the induced current $\mathbf{J}$. Using the method of moments (Peterson et al., 1998), the total field $\mathbf{E}^{\text{t}}$ for a given transmitter can be expressed as (Colton and Kress, 2013):
@@
Figure 2: Difficulties that previous methods faced under a single-transmitter setting. (a) BP cannot reconstruct the scatterer. (b) Physics-Net makes incorrect guesses. (c) Although the reconstuction result of Img-Interiors is consistent with the measured field, the reconstructed scatterer itself is completely different from the ground truth.
Reduction of measured data. EISP is fundamentally challenged by nonlinearity and ill-posedness, especially when the amount of measured data is significantly reduced, such as under single-transmitter settings. We divide previous work into three categories and systematically analyze the difficulties they faced under this setting. (a) Conventional numerical approaches, such as BP (Belkebir et al., 2005), employ linear approximations, which limit their reconstruction quality. As shown in Fig.˜2, BP cannot even reconstruct a rough shape of the scatterer. (b) Machine learning methods based on conventional numerical approaches Song et al. (2021); Liu et al. (2022), such as Physics-Net. Although Physics-Net can leverage data-driven training to compensate for missing physical information, its strong dependency on BP initialization becomes a critical bottleneck. When BP fails, the model cannot correct the error of BP because it is not fully end-to-end, resulting inaccurate reconstructions, as shown in Fig.˜2. (c) Machine learning methods based on implicit functions, such as Img-Interiors (Luo et al., 2024). Img-Interiors reconstructs a scatterer through case-by-case optimization. As shown in Fig.˜2, we use the scatterer reconstructed by Img-Interiors to simulate the scattered field, and the result closely matches the measured field. However, the scatterer itself deviates significantly from the ground truth, which shows the intrinsic ambiguity of the inverse problem. The core conclusion is that the severe information deficit makes a direct solution to the inverse problem fundamentally intractable. Consequently, any such attempt is bound to be fragile, highlighting the need for an alternative paradigm.
@@ 4.1 Overview
To address the aforementioned limitations, we introduce our end-to-end, data-driven framework for EISP, as illustrated in Fig.˜3. Our method employs an MLP that takes space coordinates $\bf{x}$} and corresponding scattered field measurements $\mathbf{E}^{\text{s}}$ as input, and directly outputs the relative permittivity $\bm{\epsilon}_{r}$ at the specified locations. This approach effectively learns the mapping between scattered field $\mathbf{E}^{\text{s}}$ and relative permittivity $\bm{\epsilon}_{r}$ through training on diverse scattering scenarios, thereby incorporating essential data distribution priors to compensate for the lack of physical information caused by insufficient measurements.
In the following, we detail our model architecture (Section˜4.2), and the training losses (Section˜4.3).
@@ 4.2 Model Architecture
where $\mathbf{x}_{\text{i}}$ represents the spatial coordinate, $\mathbf{E}^{\text{s}}$ denotes the scattered field measured by all receivers, $F_{\theta}(\cdot)$ is an MLP with trainable parameters, and $\hat{\bm{\epsilon}_{r}}(\mathbf{x}_{\text{i}})$ is the predicted relative permittivity at the corresponding position. Recall that in Sec.˜3, for a single transmitter, the scattered field $\mathbf{E}^{\text{s}}$ is discretized as a real-valued vector of dimension $2N_{r}$, containing the real and imaginary parts of the measurements from all $N_{r}$ receivers. In the multiple transmitter configuration, $\mathbf{E}^{\text{s}}$ is constructed by combining the complex measurement data from all $N$ transmitters, resulting in a real-valued vector of dimension $2N\cdot N_{r}$ that represents the wave propagation and scattering behavior under diverse illumination conditions provided by transmitters at different locations.
Figure 3: Overview of our method. Our pipeline is built around a MLP that serves as the inverse solver. Given the scattered field measurements $\mathbf{E}^{\text{s}}$ from all transmitters and receivers, along with a spatial query $\bf{x}$}, the MLP directly predicts thee relative permittivity $\hat{\bm{\epsilon}_{r}}({\bf{x}})$. To enhance spatial expressiveness, we apply positional encoding $\gamma({\bf{x}})$ to the query position. During training, dashed lines indicate the supervision signals applied.
To enhance the model’s capacity to represent high-frequency features, we apply positional encoding to the spatial coordinates $\mathbf{x}_{\text{i}}$, mapping them into a higher-dimensional Fourier feature space using the encoding function: $\gamma(x)=[\sin(x),\cos(x),\ldots,\sin(2^{\Omega-1}x),\cos(2^{\Omega-1}x)]$, where the hyperparameter $\Omega$ controls the spectral bandwidth.
The complete relative permittivity distribution $\hat{\bm{\epsilon}_{r}}$ is reconstructed by sampling the MLP at all grid points $\{\mathbf{x}_{\text{i}}\}_{i=1}^{M^{2}}$: $\hat{\bm{\epsilon}_{r}}=\{F_{\theta}(\mathbf{E}^{\text{s}},\gamma(\mathbf{x}_{\text{i}})\}_{\text{i}=1}^{M^{2}}.$

### v6

@@ Abstract
Electromagnetic Inverse Scattering Problems (EISP) seek to reconstruct relative permittivity from scattered fields and are fundamental to applications like medical imaging. This inverse process is inherently ill-posed and highly nonlinear, making it particularly challenging, especially under sparse transmitter setups, e.g., with only one transmitter. While recent machine learning-based approaches have shown promising results, they often rely on time-consuming, case-specific optimization and perform poorly under sparse transmitter setups. To address these limitations, we revisit EISP from a data-driven perspective. The scarcity of transmitters leads to an insufficient amount of measured data, which fails to capture adequate physical information for stable inversion. Accordingly, we propose a fully end-to-end and data-driven framework that predicts the relative permittivity of scatterers from measured fields, leveraging data distribution priors to compensate for the incomplete information from sparse measurements. This design enables data-driven training and feed-forward prediction of relative permittivity while maintaining strong robustness to transmitter sparsity. Extensive experiments show that our method outperforms state-of-the-art approaches in reconstruction accuracy and robustness. Notably, we demonstrate, for the first time, high-quality reconstruction from a single transmitter. This work advances practical electromagnetic imaging by providing a new, cost-effective paradigm to inverse scattering. Code and models are released at https://gomenei.github.io/SingleTX-EISP/.
Figure 1: Comparison between our method and the previous state-of-the-art. Left: Img-Interiors [27] requires case-specific optimization to reconstruct the permittivity. In contrast, our method is a data-driven framework that operates in an end-to-end, feed-forward manner for solving inverse scattering. Right: Our method yields more accurate reconstructions than Img-Interiors [27]. It remains robust even with a single transmitter and achieves real-time inference with over $70,000\times$ speed-up.
@@ 1 Introduction
Electromagnetic waves can penetrate object surfaces, making them essential for non-invasive imaging [17, 30]. At the core of electromagnetic imaging lies the Electromagnetic Inverse Scattering Problems (EISP), which seeks to reconstruct an object’s relative permittivity from measured scattered electromagnetic field [29]. By solving EISP, we can accurately recover internal structures without physical intrusion [38], enabling a range of scientific and industrial applications, such as safer and more cost-effective alternatives to X-rays and MRI scans [5, 30, 29]. Typically, EISP necessitate a large number of transmitters and receivers to acquire sufficient measurement data. This requirement, however, leads to increased operational time and higher costs, thereby limiting the practical applicability of electromagnetic imaging techniques [22]. In contrast, reducing the number of transmitters offers significant advantages, including lower costs and easier deployment in constrained environments.[3, 1, 49]
@@
However, the inherent ill-posed nature of EISP poses significant challenges to accurate reconstruction [32, 7, 23, 56, 27, 51, 50, 52, 42], particularly when only a limited number of transmitters are available. The scarcity of transmitters leads to an insufficient amount of measured data, which fails to capture adequate physical information for stable inversion. As a result, approaches relying solely on physical mechanisms[37, 4, 6, 55] often fail to achieve accurate reconstruction. Conventional numerical methods such as backpropagation (BP) [4], generally fail to produce reliable reconstructions under such limited-data conditions. Recent machine learning-based approaches like PGAN [39] and Physics-Net [26] often start with an initial solution derived from numerical methods, i.e., BP, and frame the problem as an image-to-image translation task. With only a limited number of transmitters available, reliance on BP becomes a critical bottleneck. When BP fails, these methods are unable to correct its errors, as they are not fully end-to-end, ultimately leading to inaccurate reconstructions. The most recent method Img-Interiors [27] integrates physical mechanisms into neural networks and performs case-by-case optimization. However, in limited-transmitter scenarios, even after optimization has converged, the resulting reconstructions may still diverge substantially from the ground truth (Fig.˜2), underscoring the intrinsic ambiguity of the inverse problem.
@@
To address these limitations, we propose a fully end-to-end and data-driven framework that predicts the relative permittivity of scatterers from measured fields, leveraging data distribution priors to compensate for the scarcity of observational data. Unlike generative methods where data distribution priors typically refer to explicit, decoupled modules (e.g., denoisers modeling $p(x)$) [21, 10], our method learns an end-to-end mapping that implicitly leverages data statistics to resolve the inherent ill-posedness of electromagnetic inverse scattering problems. Specifically, our model takes the measured fields and the spatial coordinate of a position as input and directly predicts the relative permittivity at that location using Multilayer Perceptron (MLP)s, and is trained in a fully end-to-end manner against the ground-truth data. Our approach bypasses traditional numerical methods like BP, thereby avoiding the inherent constraints associated with conventional inversion techniques in limited-transmitter scenarios and fully exploiting the advantages of data-driven learning. This simple yet effective design enables efficient training across datasets and supports fast, feed-forward inference to achieve accurate and stable reconstruction predictions.
@@ 2 Related Work
2.1 Electromagnetic Inverse Scattering Problems
Solving EISP is to determine the relative permittivity of the scatterers based on the scattered field measured by the receivers, thereby obtaining internal imaging of the object. The primary challenges of EISP arise from its nonlinearity, ill-posedness, and errors introduced by the discretization [32, 7, 23, 56, 27]. Traditional methods for solving EISP can be categorized into non-iterative [37, 13, 18, 4] and iterative [6, 55, 48, 19, 41] approaches. Non-iterative methods, such as the Born approximation [37], the Rytov approximation [13, 18], and the BP method [4], solve nonlinear equations through linear approximations, which inevitably lead to poor quality of the results. For better reconstruction quality, iterative methods [54, 6, 55, 48, 19, 41, 14] such as 2-fold Subspace Optimization Method (SOM) [54] and Gs SOM [6] are proposed. To further overcome the ill-posedness of EISP, diverse regularization approaches and prior information have been widely applied [31, 36, 25, 2]. However, all of these methods are not generalizable and can be time-consuming because of the iterative schemes [26].
Figure 2: Difficulties that previous methods faced under a single-transmitter setting. (a) BP cannot reconstruct the scatterer. (b) Physics-Net makes incorrect guesses. (c) Although the reconstruction result of Img-Interiors is consistent with the measured field, the reconstructed scatterer itself is completely different from the ground truth.
@@ 2.2 Machine Learning for EISP
Recent studies shift to leveraging neural networks to solve this problem and demonstrate promising results [16, 24]. Some work [45, 23, 53, 47, 35, 39, 26] adopt a two-stage strategy: they use non-iterative methods such as BP [4] to generate initial estimates, which are then refined using image-to-image neural networks. While these approaches offer a degree of generalization, they are not end-to-end and remain dependent on BP initialization [4], which becomes their bottleneck. When physical data are too insufficient to reconstruct the scatterer, especially under single-transmitter settings, these approaches tend to “hallucinate” outputs according to unreliable initialization rather than predict the scatterer based on measured field (see Fig.˜5). A more recent approach, Img-Interiors [27], integrates scattering mechanisms into the network architecture and achieves accurate reconstructions. However, it requires case-specific optimization, limiting generalization and making it vulnerable to local minima, often leading to failure in complex settings (see Figs.˜4 and 7). Moreover, it fails under a single transmitter setting even when the optimization may have already converged because of ambiguity. While our method is also learning-based, it is an end-to-end feed-forward framework that simultaneously achieves generalization through data-driven learning. As a result, it consistently outperforms SOTA methods, particularly in the challenging single-transmitter setting where previous approaches fail.
@@ 3 Revisiting EISP
Preliminary. In the forward process, the transmitters produce incident electromagnetic field $\mathbf{E}^{\text{i}}$ to the scatterer, generating scattered electromagnetic field $\mathbf{E}^{\text{s}}$. EISP is the inverse problem of the forward process. That is, for an unknown scatterer, we use transmitters to apply certain incident field $\mathbf{E}^{\text{i}}$ to it, and measure the scattered field $\mathbf{E}^{\text{s}}$ as our input via receivers. Our goal is to reconstruct the relative permittivity $\bm{\epsilon}_{r}$ throughout the scatterer. For a detailed introduction of EISP, please refer to our supplementary material (Appx.˜B). Specifically, the incident field $\mathbf{E}^{\text{i}}$ excites the induced current $\mathbf{J}$. Using the method of moments [33], the total field $\mathbf{E}^{\text{t}}$ for a given transmitter can be expressed as [8]:
@@
Figure 3: Overview of our method. Our pipeline is built around a MLP that serves as the inverse solver. Given the scattered field measurements $\mathbf{E}^{\text{s}}$ from all transmitters and receivers, along with a spatial query $\bf{x}$}, the MLP directly predicts the relative permittivity $\hat{\bm{\epsilon}_{r}}({\bf{x}})$. To enhance spatial expressiveness, we apply positional encoding $\gamma({\bf{x}})$ to the query position. During training, dashed lines indicate the supervision signals applied.
Reduction of measured data. EISP is fundamentally challenged by nonlinearity and ill-posedness, especially when the amount of measured data is significantly reduced, such as under single-transmitter settings. We divide previous work into three categories and systematically analyze the difficulties they faced under this setting. (a) Conventional numerical approaches, such as BP [4], employ linear approximations, which limit their reconstruction quality. As shown in Fig.˜2, BP cannot even reconstruct a rough shape of the scatterer. (b) Machine learning methods based on conventional numerical approaches [39, 26], such as Physics-Net. Although Physics-Net can leverage data-driven training to compensate for missing physical information, its strong dependency on BP initialization becomes a critical bottleneck. When BP fails, the model cannot correct the error of BP because it is not fully end-to-end, resulting inaccurate reconstructions, as shown in Fig.˜2. (c) Machine learning methods based on implicit functions, such as Img-Interiors [27]. Img-Interiors reconstructs a scatterer through case-by-case optimization. As shown in Fig.˜2, we use the scatterer reconstructed by Img-Interiors to simulate the scattered field, and the result closely matches the measured field. However, the scatterer itself deviates significantly from the ground truth, which shows the intrinsic ambiguity of the inverse problem. The core conclusion is that the severe information deficit makes a direct solution to the inverse problem fundamentally intractable. Consequently, any such attempt is bound to be fragile, highlighting the need for an alternative paradigm.
@@ 4.1 Overview
To address the aforementioned limitations, we introduce our end-to-end, data-driven framework for EISP, as illustrated in Fig.˜3. Our method employs an MLP that takes space coordinates $\bf{x}$} and corresponding scattered field measurements $\mathbf{E}^{\text{s}}$ as input, and directly outputs the relative permittivity $\bm{\epsilon}_{r}$ at the specified locations. This approach effectively learns the mapping between scattered field $\mathbf{E}^{\text{s}}$ and relative permittivity $\bm{\epsilon}_{r}$ through training on diverse scattering scenarios, thereby incorporating essential data distribution priors to compensate for the lack of physical information caused by insufficient measurements. In the following, we detail our model architecture (Sec.˜4.2), and the training losses (Sec.˜4.3).
@@ 4.2 Model Architecture
where $\mathbf{x}_{\text{i}}$ represents the spatial coordinate, $\mathbf{E}^{\text{s}}$ denotes the scattered field measured by all receivers, $F_{\theta}(\cdot)$ is an MLP with trainable parameters, and $\hat{\bm{\epsilon}_{r}}(\mathbf{x}_{\text{i}})$ is the predicted relative permittivity at the corresponding position. Recall that in Sec.˜3, for a single transmitter, the scattered field $\mathbf{E}^{\text{s}}$ is discretized as a real-valued vector of dimension $2N_{r}$, containing the real and imaginary parts of the measurements from all $N_{r}$ receivers. In the multiple transmitter configuration, $\mathbf{E}^{\text{s}}$ is constructed by combining the complex measurement data from all $N$ transmitters, resulting in a real-valued vector of dimension $2N\cdot N_{r}$ that represents the wave propagation and scattering behavior under diverse illumination conditions provided by transmitters at different locations. To enhance the model’s capacity to represent high-frequency features, we apply positional encoding to the spatial coordinates $\mathbf{x}_{\text{i}}$, mapping them into a higher-dimensional Fourier feature space using the encoding function: $\gamma(x)=[\sin(x),\cos(x),\ldots,\sin(2^{\Omega-1}x),\cos(2^{\Omega-1}x)]$, where the hyperparameter $\Omega$ controls the spectral bandwidth. The complete relative permittivity distribution $\hat{\bm{\epsilon}_{r}}$ is reconstructed by sampling the MLP at all grid points $\{\mathbf{x}_{\text{i}}\}_{i=1}^{M^{2}}$: $\hat{\bm{\epsilon}_{r}}=\{F_{\theta}(\mathbf{E}^{\text{s}},\gamma(\mathbf{x}_{\text{i}}))\}_{\text{i}=1}^{M^{2}}.$